OpenStudy (anonymous):

Check answers? Can someone please help me ASAP?! A student is assessing the correlation between the average number of hours of playing video games in a week and average score on a math test for the students of a class. The table below shows the data: Number of hours of playing video games (x) 0 0.5 1 1.5 2 2.5 3 3.5 4 4.5 Score on math test (y) 90 88 86 84 82 80 78 76 74 72 Part A: Write a function which best fits the data. (3 points) Part B: What does the slope and y-intercept of the plot indicate? (3 points)

OpenStudy (anonymous):

Would parta be f(x) = 90-4x?

jimthompson5910 (jim_thompson5910):

you are correct

jimthompson5910 (jim_thompson5910):

you can also write that as f(x) = -4x + 90

OpenStudy (anonymous):

Thank you! And can I have some help with part B? Im a little confused, I know the slope is -4 but I dont understand the question

jimthompson5910 (jim_thompson5910):

the slope of -4 means that as x increases by 1, y decreases by 4 slope = (change in y)/(change in x) = (-4)/(1) = -4

jimthompson5910 (jim_thompson5910):

in this specific context, the slope means "for every additional hour of playing video games, the test score decreases by 4"

jimthompson5910 (jim_thompson5910):

what does the y-intercept represent?

OpenStudy (anonymous):

y-intercept f(x)=90-4x 0=90-4x x=90/4

jimthompson5910 (jim_thompson5910):

that's the x-intercept

jimthompson5910 (jim_thompson5910):

to find the y-intercept you plug in x = 0 and evaluate

OpenStudy (anonymous):

Thank you for trying to explain this to me! I dont really understand it so Im going to ask my teacher for a tutoring session, thank you again

jimthompson5910 (jim_thompson5910):

y = -4x + 90 y = -4(0) + 90 ... replace x with 0 y = 0 + 90 y = 90 so the y-intercept is 90

jimthompson5910 (jim_thompson5910):

what does this mean? it means that if you don't play any video games at all, the score you get is 90. In a sense, this is like the starting value.

OpenStudy (anonymous):

Ohhh, ok, gosh...that was simple... I appreciate it so much

jimthompson5910 (jim_thompson5910):

no problem